# Upper bounds on the genus of Albanese fibrations

### v1

@@ 3. Upper bound on hyperelliptic Albanese fibrations with $p_{g}=q=1$
4. Quadratic upper bounds on hyperelliptic Albanese fibrations
The main purpose of this section is to prove Theorem 1.1. We will prove the following upper bounds (using the relative invariants) on the genus $g$ for any locally non-trivial hyperelliptic fibration, not necessarily the Albanese fibration. We remark once more that the relative invariants of $f$ equal the corresponding invariants of the surface $S$ (cf. (2.4)) if the base $B$ is an elliptic curve.
Theorem 4.1.
Let $f:\,S\to B$ be a relatively minimal hyperelliptic fibration of genus $g\geq 2$. Suppose that $f$ is locally non-trivial.
  1. (1)
    If $\lambda_{f}<4$, then $g\leq\frac{4\chi_{f}}{4\chi_{f}-K_{f}^{2}}$.
  2. (2)
    If $\lambda_{f}=4$, then $g\leq 2\chi_{f}+2$.
  3. (3)
    If $\lambda_{f}>4$, then
    $g\leq\left\{\begin{aligned} &2\chi_{f}+1,&&\text{if $g(B)=0$};\\ &\frac{25}{2}\chi_{f}^{2}+14\chi_{f}+4,&\quad&\text{if $g(B)=1$};\\ &32\chi_{f}^{2}+20\chi_{f}+3,&&\text{if $g(B)\geq 2$}.\end{aligned}\right.$
We first prove Theorem 1.1 based on Theorem 4.1.
Proof of Theorem 1.1.
Let $f:\,S\to B$ be the Albanese fibration of $S$ and $g$ the genus of a general fiber. Then $g(B)=q\geq 1$. If $g(B)\geq 2$, by the non-negativity of the Hodge bundle, one obtains
$\chi(\mathcal{O}_{S})-(g-1)(g(B)-1)=\chi_{f}\geq 0.$
Hence $g\leq\frac{\chi(\mathcal{O}_{S})}{g(B)-1}+1\leq\chi({\mathcal{O}_{S}})+1$. If $g(B)=1$, then $f$ must locally non-trivial; otherwise $\chi(\mathcal{O}_{S})=\chi_{f}=0$ which is absurd. Thus by (2.4) together with Theorem 4.1 (3), one obtains a quadratic upper bound on the genus $g$. ∎
We next come back to the proof of Theorem 4.1. The proof relays heavily on the formulas given by Xiao in Theorem 2.1.
Proof of Theorem 4.1.
The first statement follows from the slope inequality (2.1). The second one was proved by Ishida [Ish05]. Here we can give a short proof based on Theorem 2.1. In fact, since $\lambda_{f}=4$, i.e., $K_{f}^{2}=4\chi_{f}$, by Theorem 2.1 one gets
$\chi_{f}$ $\,=\frac{g}{4}\left(K_{f}^{2}-\frac{4(g-1)}{g}\chi_{f}\right)$
$\,=\frac{g-2}{2}.$
Thus $g\leq 2\chi_{f}+2$ as required.
It remains to prove (3). If $g(B)=0$, then
$\chi_{f}=\chi(\mathcal{O}_{S})+(g-1)=p_{g}-q+g\geq\frac{g-1}{2}.$
The inequality above follows from the inequality $q\leq\frac{g+1}{2}$ by Xiao [Xia87b] and the non-negativity of $p_{g}$. Hence $g\leq 2\chi_{f}+1$ as required. We consider next the case when $g(B)=1$, and we are going to prove by contradiction. Suppose that
$g>\frac{25}{2}\chi_{f}^{2}+14\chi_{f}+4.$ (4.1)
Then it is clear that $s_{g+2}=0$ by (2.6). Let
$k_{0}=\max\left\{k~{}\big{|}~{}s_{2k}\neq 0,\text{~{}or~{}}s_{2k+1}\neq 0\right\}.$
Then by (2.6) one obtains that
$k_{0}\leq\frac{5\chi_{f}+2}{2}.$ (4.2)
In fact, suppose on the contrary. Then $k_{0}\geq\frac{5\chi_{f}+3}{2}$, and hence from (2.6) it follows that
$\chi_{f}\geq\min\left\{\frac{(4k_{0}-1)g-4k_{0}^{2}}{5g+4},~{}\frac{2(k_{0}-1)(g-k_{0})}{5g+4}\right\}\geq\frac{\big{(}5\chi_{f}+1)\cdot\left(g-\frac{5\chi_{f}+3}{2}\right)}{5g+4}.$
It gives a contradiction to (4.1). Let
$r=\frac{k_{0}(k_{0}-1)}{g},\quad t=\lambda_{f}-\frac{4(g-1)}{g},\quad a_{k}=\frac{4k(g-k)-g}{g},\quad b_{k}=\frac{2k(g-k+1)-2g}{g}.$
Then by (2.2),
$\big{(}t\chi_{f}\big{)}^{2}+(3+4r)t\chi_{f}+4r^{2}+6r$ (4.3)
$=$ $\left(\sum_{k=1}^{k_{0}}a_{k}s_{2k+1}+\sum_{k=2}^{k_{0}}b_{k}s_{2k}\right)^{2}+(3+4r)\left(\sum_{k=1}^{k_{0}}a_{k}s_{2k+1}+\sum_{k=2}^{k_{0}}b_{k}s_{2k}\right)+4r^{2}+6r$
$\geq$ $\left(\sum_{k=1}^{k_{0}}a_{k}^{2}s_{2k+1}^{2}+\sum_{k=2}^{k_{0}}b_{k}^{2}s_{2k}^{2}\right)+(3+4r)\left(\sum_{k=1}^{k_{0}}a_{k}s_{2k+1}+\sum_{k=2}^{k_{0}}b_{k}s_{2k}\right)+4r^{2}+6r$
$=$ $\sum_{k=1}^{k_{0}}\big{(}a_{k}^{2}s_{2k+1}+(3+4r)a_{k}\big{)}s_{2k+1}+\sum_{k=2}^{k_{0}}\big{(}b_{k}^{2}s_{2k}+(3+4r)b_{k}\big{)}s_{2k}+4r^{2}+6r$
$\geq$ $\sum_{k=1}^{k_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}(4k^{2}-2k-2)s_{2k}.$
The last inequality above holds because
{$&a_{k}^{2}s_{2k+1}+(3+4r)a_{k}\geq 8k^{2}+4k-1,&~{}&\text{if $s_{2k+1}\neq 0$ with $1\leq k\leq k_{0}$};\\ &b_{k}^{2}s_{2k}+(3+4r)b_{k}\geq 4k^{2}-2k-2,&&\text{if $s_{2k}\neq 0$ with $2\leq k\leq k_{0}-1$};\\ &\big{(}b_{k_{0}}^{2}s_{2k_{0}}+(3+4r)b_{k_{0}}\big{)}s_{2k_{0}}+4r^{2}+6r\geq(4k_{0}^{2}-2k_{0}-2)s_{2k_{0}},&&\text{if $s_{2k_{0}}\neq 0$}.\end{aligned}\right.$
On the other hand, by (4.2) together with Theorem 2.1,
$\chi_{f}$ $\,=\frac{1}{8}s_{2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{k}{2}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{k}{4}s_{2k}-\frac{1}{2g+1}\left(\frac{1}{8}s_{2}+\sum_{k=1}^{[\frac{g}{2}]}\frac{2k^{2}+k}{2}s_{2k+1}+\sum_{k=2}^{[\frac{g+1}{2}]}\frac{2k^{2}-k}{4}s_{2k}\right)$
$\,=\frac{1}{8}s_{2}+\sum_{k=1}^{k_{0}}\frac{k}{2}s_{2k+1}+\sum_{k=2}^{k_{0}}\frac{k}{4}s_{2k}-\frac{1}{2g+1}\left(\frac{1}{8}s_{2}+\sum_{k=1}^{k_{0}}\frac{2k^{2}+k}{2}s_{2k+1}+\sum_{k=2}^{k_{0}}\frac{2k^{2}-k}{4}s_{2k}\right);$
$s_{2}$ $\,=e_{f}+2s_{g+2}-\sum_{k=1}^{[\frac{g}{2}]}s_{2k+1}-2\sum_{k=2}^{[\frac{g+1}{2}]}s_{2k}=e_{f}-\sum_{k=1}^{k_{0}}s_{2k+1}-2\sum_{k=2}^{k_{0}}s_{2k}.$
Hence
$e_{f}+\sum_{k=1}^{k_{0}}(4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}2(k-1)s_{2k}-8\chi_{f}$
=\,= $\frac{1}{2g+1}\bigg{(}e_{f}+\sum_{k=1}^{k_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}(4k^{2}-2k-2)s_{2k}\bigg{)}.$
Note that the left hand of the above equality is an integer, and that the right hand of the above equality is positive. Therefore,
$e_{f}+\sum_{k=1}^{k_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{k_{0}}(4k^{2}-2k-2)s_{2k}\geq 2g+1$
Combining with (4.3), one obtains
$\big{(}t\chi_{f}\big{)}^{2}+(3+4r)t\chi_{f}+4r^{2}+6r+(12-\lambda_{f})\chi_{f}$
=\,= $\big{(}t\chi_{f}\big{)}^{2}+(3+4r)t\chi_{f}+4r^{2}+6r+e_{f}$
@@ 4. Quadratic upper bounds on hyperelliptic Albanese fibrations
Hence
$\big{(}t\chi_{f}\big{)}^{2}+\big{(}(3+4r)t+12-\lambda_{f}\big{)}\chi_{f}+4r^{2}+6r\geq 2g+1>25\chi_{f}^{2}+28\chi_{f}+9.$ (4.4)
Note that $\lambda_{f}\leq 9$ by (2.5), and $t=\lambda_{f}-\frac{4(g-1)}{g}\leq\frac{5g+4}{g}$. Moreover, by (4.1) and (4.2),
$r=\frac{k_{0}(k_{0}-1)}{g}$ $\,\leq\frac{\frac{5\chi_{f}+2}{2}\cdot\frac{5\chi_{f}}{2}}{\frac{25}{2}\chi_{f}^{2}+14\chi_{f}+4}<\frac{1}{2};$
$(3+4r)t+12-\lambda_{f}$ $\,=(2+4r)\lambda_{f}-\frac{4(3+4r)(g-1)}{g}+12$
$\,\leq 9(2+4r)-\frac{4(3+4r)(g-1)}{g}+12=\frac{(16+20g)r+6g+12}{g}+12$
$\,<\frac{(16+20g)\cdot\frac{1}{2}+6g+12}{g}+12=28+\frac{20}{g}.$
Therefore, (4.4) implies that
$425\chi_{f}^{2}+28\chi_{f}+9<\bigg{(}\frac{5g+4}{g}\bigg{)}^{2}\chi_{f}^{2}+\bigg{(}28+\frac{20}{g}\bigg{)}\chi_{f}+$ 4 .
It follows that $5<\frac{40\chi_{f}^{2}+20\chi_{f}}{g}+\frac{16\chi_{f}^{2}}{g^{2}}$, which contradicts (4.1). Finally for the case when $g(B)\geq 2$, the proof would be the same as above. We simply argue as follows. Suppose that
$g>32\chi_{f}^{2}+20\chi_{f}+3.$ (4.5)
Then $s_{g+2}=0$ by (2.7). Let
$\tilde{k}_{0}=\max\left\{k~{}\big{|}~{}s_{2k}\neq 0,\text{~{}or~{}}s_{2k+1}\neq 0\right\}.$
Then by (2.7) one shows similarly that
$\tilde{k}_{0}\leq 4\chi_{f}+1.$ (4.6)
Let
$\tilde{r}=\frac{\tilde{k}_{0}(\tilde{k}_{0}-1)}{g},\quad t=\lambda_{f}-\frac{4(g-1)}{g},\quad a_{k}=\frac{4k(g-k)-g}{g},\quad b_{k}=\frac{2k(g-k+1)-2g}{g}.$
Then by (2.2),
$\big{(}t\chi_{f}\big{)}^{2}+(3+4\tilde{r})t\chi_{f}+4\tilde{r}^{2}+6\tilde{r}\geq\sum_{k=1}^{\tilde{k}_{0}}(8k^{2}+4k-1)s_{2k+1}+\sum_{k=2}^{\tilde{k}_{0}}(4k^{2}-2k-2)s_{2k}.$
And hence
$\big{(}t\chi_{f}\big{)}^{2}+\big{(}(3+4\tilde{r})t+12-\lambda_{f}\big{)}\chi_{f}+4\tilde{r}^{2}+6\tilde{r}$
=\,= $\big{(}t\chi_{f}\big{)}^{2}+(3+4\tilde{r})t\chi_{f}+4\tilde{r}^{2}+6\tilde{r}+e_{f}$
[…]
Similarly by applying (2.3) instead of (2.5), one obtains
$464\chi_{f}^{2}+40\chi_{f}+7<\bigg{(}\frac{8g+4}{g}\bigg{)}^{2}\chi_{f}^{2}+\bigg{(}40+\frac{20}{g}\bigg{)}\chi_{f}+$ 4 .
It follows that $3<\frac{64\chi_{f}^{2}+20\chi_{f}}{g}+\frac{16\chi_{f}^{2}}{g^{2}}$, which contradicts (4.5). This completes the proof. ∎
[…]
(1) From the proof above, one can even obtain a quadratic upper bound in Theorem 4.1 (3) such that the quadratic coefficient relies on the slope $\lambda_{f}$.
(2) The quadratic upper bounds given in Theorem 4.1 (3) seem not sharp. Nevertheless, we will construct examples in the next section showing that the quadratic upper bounds can not improved into linear ones.
In this section, we will construct examples showing that the genus $g$ of hyperelliptic Albanese fibrations can be as large as a quadratic function in $K_{S}^{2}$ or $\chi(\mathcal{O}_{S})$.
Theorem 5.1.
There exist a sequence of minimal irregular surfaces $S_{k}$ ($k\geq 3$ and odd) of general type with $q(S_{k})=1$, such that their Albanese maps induce a hyperelliptic fibration $f_{k}:\,S_{k}\to E$ of genus $g_{k}=\frac{(k-1)(k+3)}{2}$, and that
$\chi(\mathcal{O}_{S_{k}})=\chi_{f_{k}}=\frac{3}{2}(k-1),\qquad K_{S_{k}}^{2}=K_{f_{k}}^{2}=8k-12.$
@@ 5. Hyperelliptic Albanese fibrations with a quadratic Albanese genus
Let $h:P\rightarrow E$ be the $\mathbb{P}^{1}$-bundle and above. For any irreducible curve $D\subseteq P$, let $D\sim aC_{0}+b\Gamma$. Then $a\geq 0$, $b\geq 0$, and $a+b>0$. Moreover, if $b=0$, then $a\geq m$ unless $D=C_{0}$ or $C_{\infty}$.
@@
Let $h:P\rightarrow E$ be the $\mathbb{P}^{1}$-bundle and above. For any $p\in P\setminus\{C_{0}\cup C_{\infty}\}$, let $\tau:\,\widetilde{P}\to P$ be the blowing-up centered at $p$, and $\mathcal{E}$ be the exceptional curve. Then for any odd $k\geq 3$ and $g_{k}=\frac{(k-1)(k+3)}{2}$, there exists a smooth divisor $\widetilde{R}\in\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$ if $m\geq k+2$.
@@
Note that numerically we have $K_{P}\sim-2C_{0}$. Hence $L^{2}=4k+3\geq 15>0$. We claim that
@@
is base-point-free by applying Reider’s method [Rei88]. Hence a general divisor $\widetilde{R}\in\left|K_{\widetilde{P}}+L\right|$ is smooth by Bertini’s theorem. It remains to prove Section 5. Let $\widetilde{D}\subseteq\widetilde{P}$ be any irreducible curve. If $\widetilde{D}=\mathcal{E}$, or $\widetilde{D}$ is the strict transform of any fiber of $h:\,P\to E$, then one checks easily that
@@
$L\cdot\widetilde{D}=2\big{(}a+(g_{k}+2)b\big{)}-(2k+1)\beta=2a+(k+1)^{2}b-(2k+1)\beta.$
If $\beta>k$, i.e., $\beta\geq k+1$, then
$L\cdot\widetilde{D}$ $\,=2+2(a-1)+(k+1)^{2}b-(2k+1)\beta$
@@
$\,>2+2(k+1)\big{(}\beta-\frac{1}{2}\big{)}-(2k+1)\beta$
$\,=1+\beta-k\geq 2;$
if $\beta\leq k$, then
$L\cdot\widetilde{D}$ $\,=2+2(a-1)+(k+1)^{2}b-(2k+1)\beta$
{$&2+2(k+1)^{2}-(2k+1)k=3k+4>2,&\quad&\text{if~{}}b\geq 2;\\ &2+\beta(\beta-1)+(k+1)^{2}-(2k+1)\beta=2+(k+1-\beta)^{2}>2,&&\text{if~{}}b=1,\\ &2,&&\text{if $b=0$ and $a=1$};\\ &2a-(2k+1)\beta\geq 2m-(2k+1)>2,&&\text{if $b=0$ and $a\geq 2$};\end{aligned}\right.$
@@
We can now construct the required hyperelliptic Albanese fibrations.
Proof of Theorem 5.1.
For any odd $k\geq 3$, let $\widetilde{R}\in\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$ be any smooth divisor on $\widetilde{P}$ as in Section 5. Let $R=\tau(\widetilde{R})\subseteq P$ be its image in $P$. Then $R$ admits a singularity of order $2k$ at the point $p$, and
@@
Hence one can construct a hyperelliptic fibration $f_{k}:\,S_{k}\to E$ using the above data $(P,R,\delta)$. The genus $g_{k}$ of a general fiber of $f_{k}$ is $g_{k}=\frac{(k-1)(k+3)}{2}$. Moreover, using the notations introduced in Section 2.2, one computes that $s_{2k}=1$, $s_{2}=(K_{\widetilde{P}/E}+\widetilde{R})\cdot\widetilde{R}=10k-8$, and $s_{i}=0$ for other $i$. Hence by Theorem 2.1,
$\chi(\mathcal{O}_{S_{k}})=\chi_{f_{k}}=\frac{3}{2}(k-1),\qquad K_{S_{k}}^{2}=K_{f_{k}}^{2}=8k-12.$

### v3

@@ 5. Hyperelliptic Albanese fibrations with a quadratic Albanese genus
In this section, we will construct several examples. Section 5 and Section 5 show that the equalities in Section 3 can be reached for both $g$ odd and $g$ even; in particular, the genus $g$ of hyperelliptic Albanese fibrations can be as large as a quadratic function in $\chi(\mathcal{O}_{S})$. Section 5 indicates that the equality in Theorem 3.1 (1) is also sharp.
@@
There exist a sequence of minimal irregular surfaces $S_{k}$ ($k\geq 3$ and odd) of general type with $q(S_{k})=1$, such that their Albanese maps induce a hyperelliptic fibration $f_{k}:\,S_{k}\to E$ of odd genus $g_{k}=\frac{(k-1)(k+3)}{2}+1$, and that
$\chi(\mathcal{O}_{S_{k}})=\chi_{f_{k}}=\frac{3k-1}{2},\qquad K_{S_{k}}^{2}=K_{f_{k}}^{2}=8k-8.$
@@
Let $h:P\rightarrow E$ be the $\mathbb{P}^{1}$-bundle above. For any irreducible curve $D\subseteq P$, let $D\sim aC_{0}+b\Gamma$. Then $a\geq 0$, $b\geq 0$, and $a+b>0$. Moreover, if $b=0$, then $a\geq m$ unless $D=C_{0}$ or $C_{\infty}$.
@@
Let $h:P\rightarrow E$ be the $\mathbb{P}^{1}$-bundle above. For any $p\in P\setminus\{C_{0}\cup C_{\infty}\}$, let $\tau:\,\widetilde{P}\to P$ be the blow-up centered at $p$, and $\mathcal{E}$ be the exceptional curve. Then for any odd $23\leq k\leq m-$ 2, there exists a smooth divisor $\widetilde{R}\in\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$, where $g_{k}=\frac{(k-1)(k+3)}{2}+1$.
@@
Note that numerically we have $K_{P}\sim-2C_{0}$. Hence $L^{2}=4k+11\geq 15>0$. We claim that
@@
is base-point-free by applying Reider’s method [Rei88]. Hence a general divisor $\widetilde{R}\in\left|K_{\widetilde{P}}+L\right|$ is smooth by Bertini’s theorem.
It remains to prove Section 5. Let $\widetilde{D}\subseteq\widetilde{P}$ be any irreducible curve. If $\widetilde{D}=\mathcal{E}$, or $\widetilde{D}$ is the strict transform of any fiber of $h:\,P\to E$, then one checks easily that
@@
$L\cdot\widetilde{D}=2\big{(}a+(g_{k}+2)b\big{)}-(2k+1)\beta=2a+(k+1)^{2}b-(2k+1)\beta+2b.$
If $\beta\geq k+2$, then
$L\cdot\widetilde{D}$ $\,=2+2(a-1)+(k+1)^{2}b-(2k+1)\beta+2b$
@@
$\,>2;$
if $\beta\leq k+1$, then
$L\cdot\widetilde{D}$ $\,=2+2(a-1)+(k+1)^{2}b-(2k+1)\beta+2b$
{$&2+2(k+1)^{2}-(2k+1)(k+1)=k+3>2,&\quad&\text{if~{}}b\geq 2;\\ &2+\beta(\beta-1)+(k+1)^{2}-(2k+1)\beta=2+(k+1-\beta)^{2}>2,&&\text{if~{}}b=1,\\ &2,&&\text{if $b=0$ and $a=1$};\\ &2a-(2k+1)\beta\geq 2m-(2k+1)>2,&&\text{if $b=0$ and $a\geq 2$};\end{aligned}\right.$
@@
We come back to the construction of the required hyperelliptic Albanese fibrations in Section 5. For any odd $k\geq 3$, let $\widetilde{R}\in\left|\tau^{*}\big{(}(2g_{k}+2)C_{0}+2\Gamma\big{)}-2k\mathcal{E}\right|$ be any smooth divisor on $\widetilde{P}$ as in Section 5. Let $R=\tau(\widetilde{R})\subseteq P$ be its image in $P$. Then $R$ admits a singularity of order $2k$ at the point $p$, and
@@
Hence one can construct a hyperelliptic fibration $f_{k}:\,S_{k}\to E$ using the above data $(P,R,\delta)$. The genus $g_{k}$ of a general fiber of $f_{k}$ is $g_{k}=\frac{(k-1)(k+3)}{2}+1$. Moreover, using the notations introduced in Section 2.2, one computes that $s_{2k}=1$, $s_{2}=(K_{\widetilde{P}/E}+\widetilde{R})\cdot\widetilde{R}=10k$, and $s_{i}=0$ for other $i$. Hence by Theorem 2.1 (one can also compute $\chi(\mathcal{O}_{S_{k}})$ and $K_{S_{k}}^{2}$ by applying the formulas [BHPV04, § V.22] for double covers, as $\widetilde{R}$ is already smooth),
$\chi(\mathcal{O}_{S_{k}})=\chi_{f_{k}}=\frac{3k-1}{2},\qquad K_{S_{k}}^{2}=K_{f_{k}}^{2}=8k-8.$